# Flow-GRPO: Training Flow Matching Models via Online RL

### v4

@@ Abstract
We propose Flow-GRPO, the first method to integrate online reinforcement learning (RL) into flow matching models. Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original number of inference steps, significantly improving sampling efficiency without sacrificing performance. Empirically, Flow-GRPO is effective across multiple text-to-image tasks. For compositional generation, RL-tuned SD3.5-M generates nearly perfect object counts, spatial relations, and fine-grained attributes, increasing GenEval accuracy from $63\%$ to $95\%$. In visual text rendering, accuracy improves from $59\%$ to $92\%$, greatly enhancing text generation. Flow-GRPO also achieves substantial gains in human preference alignment. Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT-4o. (b) Image quality metrics on DrawBench [1] remain essentially unchanged. (c) Human Preference Scores on DrawBench improves after training. Results show that Flow-GRPO enhances the desired capability while preserving image quality and exhibiting minimal reward-hacking.
@@ 2 Related Work
Online RL has effectively improved the reasoning abilities of LLMs, such as DeepSeek-R1 [10] and OpenAI-o1 [11], using policy gradient methods like PPO [20] or value-free GRPO [16]. GRPO is more memory efficient by removing the need for a value network, so we adopt it in this work. PPO can also be applied to flow matching in a similar way.
@@
Recent efforts to align pretrained T2I models with human preferences follow five main directions: (1) direct fine-tuning with scalar rewards [30, 31, 32, 33]; (2) Reward Weighted Regression (RWR) [34, 35, 36]; (3) Direct Preference Optimization (DPO) and variants [37, 38, 39, 14, 40, 41, 42, 43, 44, 36]; (4) PPO-style policy gradients [45, 46, 47, 48, 49, 50]; (5) training-free alignment methods [51, 52, 53]. These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency. Building on this progress, we introduce GRPO for flow matching models, the backbone of today’s state-of-the-art T2I systems. Concurrent work [54] applies GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction by estimating a Gaussian distribution (predicting both the mean and variance of velocity), which requires retraining the pre-trained model. Another study [55] also explores SDE-based stochasticity but focuses on inference-time scaling.
@@ 5 Experiments
This section empirically evaluates Flow-GRPO’s ability to improve flow matching models on three tasks. (1) Composition Image Generation: This task requires precise object arrangement and attribute control. We report the results on GenEval. (2) Visual Text Rendering: a rule-based task that evaluates the accurate rendering of the text specified in the prompt. (3) Human Preference Alignment: This task aims to align T2I models with human preferences.
@@ 5.1 Experimental Setup
[…]
@@ 5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 10, 11 & 12 in Appendix C.4 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity, a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
We also compare Flow-GRPO with other alignment methods, including supervised fine-tuning, reward weighted regression [14, 69], Flow-DPO [14], and their online variants. Flow-GRPO outperforms all of them by a large margin. See Appendix C.1 for details.
@@ 5.3 Analysis
Figure 5 (a) highlights Denoising Reduction’s significant impact on accelerating training. To explore how different timesteps affect optimization, these experiments are conducted without the KL constraint. Reducing data collection timesteps from $40$ to $10$ achieves over a 4×4$\times$4 × speedup across all three tasks, without impacting final reward. Further reducing to $5$ does not consistently improve speed and sometimes slows training, so we choose $10$ timesteps for later experiments. For the other two tasks, learning curves of reward versus training time are presented in Figure 8 in the Appendix C.2.
@@
Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4). Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes. It also effectively controls object count, generalizing from training on $2-4$ objects to generate $5-6$ objects. Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++ [6, 70]. This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
@@
Table 4: Flow-GRPO demonstrates strong generalization. Unseen Objects: Trained on 60 object classes, evaluated on 20 unseen classes. Unseen Counting: Trained to render 2, 3, or 4 objects, was evaluated on rendering 5 or 6.
@@
Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding Counting
SD3.5-M 0.64 0.96 0.73 0.53 0.87 0.26 0.47 0.13
SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48
@@ 6 Conclusion
We have presented Flow-GRPO, the first method to integrate online RL into flow matching models. By converting deterministic ODEs to SDEs and reducing denoising steps during training, Flow-GRPO enables efficient RL-based optimization without noticeably compromising image quality or diversity. Our method significantly improves performance on compositional generation, text rendering, and human preference alignment, with minimal reward hacking. Flow-GRPO offers a simple and general framework for applying online RL to flow-based generative models.
@@ Acknowledgements
We would like to thank Mingwu Zheng for discussions on the proof and Zhanhui Zhou for suggestions on the writing.
@@ Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
$-\nabla\cdot[f_{\text{SDE}}\,p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (14)
@@ B.1 Quality Metrics
  • Aesthetic score [57]: a CLIP-based linear regressor that predicts an image’s aesthetic score.
  • DeQA score [58]: a multimodal large language model based image-quality assessment (IQA) model that quantifies how distortions, texture damage, and other low-level artefacts affect perceived quality.
@@
  • UnifiedReward [59]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state-of-the-art performance on the human preference assessment leaderboard.
@@ Appendix C Extended Experimental Results
  • Flow-DPO[14]: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps. As shown in Figure 6, Flow-GRPO outperforms all other methods. The figure also indicates that DPO and SFT improve over time. In contrast, RWR does not, which aligns with experimental findings on RWR in [12]. Additionally, Online DPO surpasses offline DPO, aligning with [15]’s finding that online DPO performs better. For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse. Figure 7 shows similar conclusions.
Figure 6: Comparison of Flow-GRPO and Other Alignment Methods on the Human Preference Alignment task.
Figure 7: Comparison of Flow-GRPO and Other Alignment Methods on the Compositional Image Generation task.
@@ C.2 Effect of Denoising Reduction
C.3 Learning Curves with or without KL
@@
C.4 Additional Qualitative Results
Figures 10, 11 & 12 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively. Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking. Conversely, removing the KL constraint significantly degrades image quality and diversity.
@@ C.4 Additional Qualitative Results
C.5 Evolution of Evaluation Images During Flow-GRPO Training

### v5

@@ Abstract
We propose Flow-GRPO, the first method to integrate online policy gradient reinforcement learning (RL) into flow matching models. Our approach uses two key strategies: (1) an ODE-to-SDE conversion that transforms a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) that matches the original model’s marginal distribution at all timesteps, enabling statistical sampling for RL exploration; and (2) a Denoising Reduction strategy that reduces training denoising steps while retaining the original number of inference steps, significantly improving sampling efficiency without sacrificing performance. Empirically, Flow-GRPO is effective across multiple text-to-image tasks. For compositional generation, RL-tuned SD3.5-M generates nearly perfect object counts, spatial relations, and fine-grained attributes, increasing GenEval accuracy from $63\%$ to $95\%$. In visual text rendering, accuracy improves from $59\%$ to $92\%$, greatly enhancing text generation. Flow-GRPO also achieves substantial gains in human preference alignment. Notably, very little reward hacking occurred, meaning rewards did not increase at the cost of appreciable image quality or diversity degradation.
Figure 1: (a) GenEval performance rises steadily throughout Flow-GRPO’s training and outperforms GPT‑4o. (b) Image quality metrics on DrawBench [1] remain essentially unchanged. (c) Human Preference Scores on DrawBench improves after training. Results show that Flow-GRPO enhances the desired capability while preserving image quality and exhibiting minimal reward-hacking.
@@ 2 Related Work
Online RL has effectively improved the reasoning abilities of LLMs, such as DeepSeek‑R1 [10] and OpenAI‑o1 [11], using policy gradient methods like PPO [20] or value-free GRPO [16]. GRPO is more memory efficient by removing the need for a value network, so we adopt it in this work. PPO can also be applied to flow matching in a similar way.
@@
Recent efforts to align pretrained T2I models with human preferences follow five main directions: (1) direct fine-tuning with differentiable rewards [30, 31, 32, 33]; (2) Reward Weighted Regression (RWR) [34, 35, 36, 37]; (3) Direct Preference Optimization (DPO) and variants [38, 39, 14, 40, 41, 42, 43, 44, 45, 46]; (4) PPO-style policy gradients [47, 48, 49, 50, 51, 52]; (5) training-free alignment methods [53, 54, 55]. These methods have successfully aligned T2I models with human preferences, improving aesthetics and semantic consistency. Building on this progress, we introduce GRPO for flow matching models, the backbone of today’s state-of-the-art T2I systems. Concurrent work [56] applies GRPO to text-to-speech flow models, but instead of converting the ODE to an SDE to inject stochasticity, they reformulate velocity prediction by estimating a Gaussian distribution (predicting both the mean and variance of velocity), which requires retraining the pre-trained model. Another study [57] also explores SDE-based stochasticity but focuses on inference-time scaling.
@@ 5 Experiments
This section empirically evaluates Flow-GRPO’s ability to improve flow matching models on three tasks. (1) Composition Image Generation: This task requires precise object arrangement and attribute control. We report the results on GenEval. (2) Visual Text Rendering: a rule‑based task that evaluates the accurate rendering of the text specified in the prompt. (3) Human Preference Alignment: This task aims to align T2I models with human preferences.
@@ 5.1 Experimental Setup
GenEval [17] assesses T2I models on complex compositional prompts—like object counting, spatial relations, and attribute binding—across six difficult compositional image generation tasks. We use its official evaluation pipeline, which detects object bounding boxes and colors, then infers their spatial relations. Training prompts are generated using official GenEval scripts, which apply templates and random combinations to construct the prompt dataset. The test set is strictly deduplicated: prompts differing only in object order (e.g., "a photo of A and B" vs. "a photo of B and A") are treated as identical, and these variants are removed from the training set. Based on the base model’s initial accuracy across the six tasks, we set the prompt ratio as $\texttt{Position}:\texttt{Counting}:\texttt{Attribute Binding}:\texttt{Colors}:\texttt{Two Objects}:\texttt{Single Object}=7:5:3:1:1:0$. Rewards are rule-based: (1) Counting: $r=1-|N_{\text{gen}}-N_{\text{ref}}|/N_{\text{ref}}$; (2) Position / Color: If the object count is correct, a partial reward is assigned; the remainder is granted when the predicted position or color is also correct.
@@
Figure 4: Comparison with Other Alignment Methods on the Compositional Generation Task.
Figure 5: Ablation Studies on Different Group Size $G$. Higher group size performs better.
@@ 5.2 Main Results
Figure 1 and Table 1 show Flow-GRPO’s GenEval performance steadily improving during training, ultimately outperforming GPT-4o. This occurs while maintaining both image quality metrics and preference scores on DrawBench, a benchmark with diverse and comprehensive prompts for evaluating general model capabilities. Figure 3 offers qualitative comparisons. Beyond Compositional Image Generation, Table 2 details evaluations on Visual Text Rendering and Human Preference tasks. Flow-GRPO improved text rendering ability, again without decreasing image quality metrics and preference scores on DrawBench. See Figures 13, 14 & 15 in Appendix C.6 for related qualitative examples. For the Human Preference task, image quality did not decrease without KL regularization. However, we found that omitting KL caused a collapse in visual diversity, a form of reward hacking discussed further in Section 5.3. These results demonstrate that Flow-GRPO boosts desired capabilities while causing very little degradation to image quality or visual diversity.
Flow-GRPO vs. Other Alignment Methods.
We compare Flow-GRPO with several alignment methods: supervised fine-tuning (SFT), Flow-DPO [14, 39], and their online variants. Flow-GRPO consistently outperforms all baselines by a significant margin. At each step, we generate a group of images using the same group size as in Flow-GRPO. The only difference lies in the update rule:
  • Flow-DPO: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
Offline variants use a fixed pretrained model for data collection, while online variants update their data collection models every 40 steps. As shown in Figure 5, Flow-GRPO outperforms all baselines. Online DPO also surpasses its offline counterpart, consistent with [15]. For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse. Appendix C presents more comprehensive comparisons covering additional methods and tasks.
@@ 5.3 Analysis
This section presents several analyses to better understand the behavior and robustness of Flow-GRPO. We examine issues such as reward hacking, the impact of denoising reduction and noise levels, the effect of group size, and the model’s generalization ability. We provide additional analyses in the Appendix C.
@@
Figure 7 (a) highlights Denoising Reduction’s significant impact on accelerating training. To explore how different timesteps affect optimization, these experiments are conducted without the KL constraint. Reducing data collection timesteps from $40$ to $10$ achieves over a $4\times$ speedup across all three tasks, without impacting final reward. Further reducing to $5$ does not consistently improve speed and sometimes slows training, so we choose $10$ timesteps for later experiments. For the other two tasks, learning curves of reward versus training time are presented in Figure 9 in the Appendix C.2.
@@
Effect of Group Size.
Figure 5 shows the effect of group size $G$ using PickScore as the reward function. When the group size was reduced to $G=12$ and $G=6$, training became unstable and eventually collapsed, whereas $G=24$ remained stable throughout the process. We observe that smaller group sizes produce inaccurate advantage estimates, increasing variance and leading to training collapse, a phenomenon also reported in [71, 72].
@@
Flow-GRPO demonstrates strong generalization on unseen scenarios from GenEval (Table 4). Specifically, it captures object number, color, and spatial relations, generalizing well to unseen object classes. It also effectively controls object count, generalizing from training on 242-4 objects to generate 565-6 or $12$ objects. Furthermore, Table 3 shows Flow-GRPO achieves significant gains on T2I-CompBench++ [6, 73]. This comprehensive benchmark for open-world compositional T2I generation features object classes and relationships substantially different from our model’s GenEval-style training data.
@@
Table 4: Flow-GRPO demonstrates strong generalization. Unseen Objects: Trained on 60 object classes, evaluated on 20 unseen classes. Unseen Counting: Trained to render 2, 3, or 4 objects, and evaluated in two settings: rendering 5 or 6 objects, and rendering 12 objects.
@@
Overall Single Obj. Two Obj. Counting Colors Position Attr. Binding 5-6 Objects 12 Objects
SD3.5-M 0.64 0.96 0.73 0.53 0.87 0.26 0.47 0.13 0.02
SD3.5-M+Flow-GRPO 0.90 1.00 0.94 0.86 0.97 0.84 0.77 0.48 0.12
@@ 6 Conclusion
We have presented Flow-GRPO, the first method to integrate online policy gradient RL into flow matching models. By converting deterministic ODEs to SDEs and reducing denoising steps during training, Flow-GRPO enables efficient RL-based optimization without noticeably compromising image quality or diversity. Our method significantly improves performance on compositional generation, text rendering, and human preference alignment, with minimal reward hacking. Flow-GRPO offers a simple and general framework for applying online RL to flow-based generative models.
@@ Acknowledgements
This work was partially supported by the JC STEM Lab of AI for Science and Engineering, funded by The Hong Kong Jockey Club Charities Trust, the Research Grants Council of Hong Kong (Project No. CUHK14213224). We gratefully acknowledge Mingwu Zheng for his insightful discussions on the proof and Zhanhui Zhou for his valuable comments that improved the clarity of this paper.
@@ Appendix A Mathematical Derivations for Stochastic Sampling using Flow Models
$\nabla\cdot[f_{\text{SDE}}\,p_{t}({\bm{x}})]+\frac{1}{2}\nabla^{2}[\sigma_{t}^{2}p_{t}({\bm{x}})]=-\nabla\cdot[{\bm{v}}_{t}({\bm{x}}_{t},t)p_{t}({\bm{x}})]$ (14)
@@ B.1 Quality Metrics
  • Aesthetic score [59]: a CLIP‑based linear regressor that predicts an image’s aesthetic score.
  • DeQA score [60]: a multimodal large language model based image‑quality assessment (IQA) model that quantifies how distortions, texture damage, and other low‑level artefacts affect perceived quality.
@@
  • UnifiedReward [61]: a recently proposed unified reward model for multimodal understanding and generation that currently achieves state‑of‑the‑art performance on the human preference assessment leaderboard.
@@ Appendix C Extended Experimental Results
  • Flow-DPO [14, 39]: Use the highest-reward image in each group as the chosen sample and the lowest as the rejected, then apply the DPO loss.
Offline variants use a fixed pretrained model for data collection, while online variants update their data collection model every 40 steps. As shown in Figure 8, Flow-GRPO outperforms all other methods. The figure also indicates that DPO and SFT improve over time. In contrast, RWR does not, which aligns with experimental findings on RWR in [12]. Additionally, Online DPO surpasses offline DPO, aligning with [15]’s finding that online DPO performs better. For the second-best online DPO, a hyperparameter search on its key parameter $\beta$ revealed that smaller values are not always optimal; excessively small $\beta$ values can cause training collapse.
Figure 8: Comparison of Flow-GRPO and Other Alignment Methods on the Human Preference Alignment task. Since methods like DPO use different tuned batch sizes from Flow-GRPO, we use the number of training prompts on the x-axis for a fair comparison across these methods.
DDPO.
DDPO [12] was originally developed for diffusion-based backbones, so we adapted it to flow-matching models via our ODE-to-SDE conversion. Using SD3.5-M as the base model and PickScore as the reward signal, we track the evaluation reward throughout the entire training process in Figure 8. We find that DDPO’s reward increases more slowly than Flow-GRPO’s and eventually collapses in the later stages, whereas Flow-GRPO trains stably and continues to improve consistently over time.
ReFL.
ReFL [32] directly fine-tunes diffusion models by viewing reward model scores as human preference losses and back-propagating gradients to a randomly-picked late timestep $t$. Following ImageReward [32], we back-propagate gradients to a randomly chosen late timestep $t\in[30,40]$ during denoising. Figure 8 shows that GRPO surpasses ReFL when the reward is differentiable, indicating that GRPO maintains strong performance in settings where ReFL applies. More importantly, GRPO does not require differentiable rewards, enabling direct use of state-of-the-art Vision-Language Models (VLMs) as reward providers. This offers two key advantages:
  • Sophisticated, General-Purpose Rewards: VLMs can conduct human-like evaluations through a structured reasoning process. Given a prompt, a VLM can decompose it into key criteria, reason step by step to verify each aspect in the generated image, and then provide a comprehensive overall score. This enables a single, unified reward model to handle diverse tasks, from text-to-image generation to complex instruction-based image editing.
  • Future-Proof and Cost-Free Upgrades: The field of VLMs is advancing at a breathtaking pace. By using a VLM as the reward source, our framework automatically benefits from these improvements. As VLMs become more capable, the reward model becomes stronger without any additional training data or computational cost.
ORW.
ORW [35] is an online reward-weighted regression method that guides the model to prioritize high-reward regions. Unlike KL regularization, it employs Wasserstein-2 regularization to prevent policy collapse and maintain diversity. To ensure a fair comparison, we adopt the same experimental setup as in our Human Preference Alignment task. For ORW, we set $\beta=0.5$ and $\alpha=1$ (lower values led to unstable training). The steps_per_epoch parameter, which controls how frequently the data-collecting policy is updated, was chosen from 20,40,100,400{20,40,100,400} based on best performance. Table 5 reports reward scores on the test set across training steps. Following ORW’s Table 1, we randomly sampled 50 DrawBench prompts and generated 64 images per prompt to compute CLIP and Diversity scores. As shown in Table 6, Flow-GRPO outperforms ORW on both metrics.
Table 5: Reward scores on the test set over training steps.
Method Step 0 Step 240 Step 480 Step 720 Step 960
SD3.5-M + ORW 28.79 29.05 29.15 27.58 23.05
SD3.5-M + Flow-GRPO 28.79 29.10 29.17 29.51 29.89
Table 6: Comparison of CLIP and diversity scores across different fine-tuning methods.
   Method    CLIP Score $\uparrow$    Diversity Score $\uparrow$
   SD3.5-M    27.99    0.96
   SD3.5-M + ORW    28.40    0.97
   SD3.5-M + Flow-GRPO    30.18    1.02
@@ C.2 Effect of Denoising Reduction
C.3 Effect of Initial Noise
We initialize each rollout with difference random noise to increase exploratory diversity during RL training. We perform an additioanl ablation to confirm this claim. With SD3.5-M as the base model and PickScore as the reward, we compare Flow-GRPO with different initial noise against Flow-GRPO with the same initial noise. Figure 11 shows the variant with different noise consistently achieved high rewards during the training process.
Figure 10: Effect of Initial Noise
Figure 11: Additional Results on FLUX.1-Dev
C.4 Additional Results on FLUX.1-Dev
We run Flow-GRPO on FLUX.1-Dev [5] using PickScore as the reward signal. The reward curve rises steadily throughout training without noticeable reward hacking. Figure 11 shows the reward values over the training process, and Table 7 compares FLUX.1-Dev with FLUX.1-Dev + Flow-GRPO on DrawBench.
Table 7: Comparison of FLUX.1-Dev and Flow-GRPO fine-tuned models.
Model Aesthetic DeQA ImageReward PickScore UnifiedReward
FLUX.1-Dev 5.71 4.31 0.85 22.62 3.65
FLUX.1-Dev + Flow-GRPO 6.02 4.24 1.32 23.97 3.81
C.5 Learning Curves with or without KL
@@
C.6 Additional Qualitative Results
Figures 13, 14 & 15 qualitatively compare SD3.5-M with its Flow-GRPO enhanced versions (with and without KL regularization) using GenEval, OCR and PickScore rewards, respectively. Flow-GRPO with KL regularization improves the target capability while maintaining image quality and minimizing reward-hacking. Conversely, removing the KL constraint significantly degrades image quality and diversity.
@@ C.6 Additional Qualitative Results
C.7 Evolution of Evaluation Images During Flow-GRPO Training